OpenStudy (anonymous):

explain how you would factor completely x9 –27. PLEASE

OpenStudy (anonymous):

do you know the formula for a Difference of two cubes? or does that sound familiar to you?

OpenStudy (anonymous):

it sounds familiar

OpenStudy (anonymous):

hmmmm. sounds familiar

OpenStudy (anonymous):

Thats what we are going to use to solve this problem, the formula you need to know is: \[x^3-y^3 = (x-y)(x^2+xy+y^2)\] @satellite lol

OpenStudy (anonymous):

so the cube root of 27 is 3 but what is the cube root of x^9??

OpenStudy (anonymous):

First, we need to get our problem in the right form. So lets rewrite it with some cubes in it: \[x^9-27 \Rightarrow (x^3)^3 - 3^3\]

OpenStudy (anonymous):

for the x^9, we change it to: \[(x^3)^3\]

OpenStudy (anonymous):

good thinking joe

OpenStudy (anonymous):

oh now i understand it so after that do i just put it in the formula???

OpenStudy (anonymous):

@irvine yep, the formula will take care of it now. @jimmy thank you :)

OpenStudy (anonymous):

oh heck no!

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

@elizated the answer is not \[(x^3+3)^3\]

OpenStudy (anonymous):

you need to use what joemath wrote above \[a^3-b^3=(a-b)(a^2+ab+b^2)\] making the replacement \[a=x^2,b=3\]

OpenStudy (anonymous):

@joe thanks that help @satellite thanks too because i got a lil confused at the end

OpenStudy (anonymous):

yes i see that. joemath did all the work, but you still have to make the substitutions

OpenStudy (anonymous):

yes i know how to do that. but a=x^3 or x^2??